# ReactDance: Hierarchical Representation for High-Fidelity and Coherent Long-Form Reactive Dance Generation

### v2

@@ Abstract
Reactive dance generation (RDG), the task of generating a dance conditioned on a lead dancer’s motion, holds significant promise for enhancing human-robot interaction and immersive digital entertainment. Despite progress in duet synchronization and motion-music alignment, two key challenges remain: generating fine-grained spatial interactions and ensuring long-term temporal coherence. In this work, we introduce ReactDance, a diffusion framework that operates on a novel hierarchical latent space to address these spatiotemporal challenges in RDG. First, for high-fidelity spatial expression and fine-grained control, we propose Hierarchical Finite Scalar Quantization (HFSQ). This multi-scale motion representation effectively disentangles coarse body posture from subtle limb dynamics, enabling independent and detailed control over both aspects through a layered guidance mechanism. Second, to efficiently generate long sequences with high temporal coherence, we propose Blockwise Local Context (BLC), a non-autoregressive sampling strategy. Departing from slow, frame-by-frame generation, BLC partitions the sequence into blocks and synthesizes them in parallel via periodic causal masking and positional encodings. Coherence across these blocks is ensured by a dense sliding-window training approach that enriches the representation with local temporal context. Extensive experiments show that ReactDance substantially outperforms state-of-the-art methods in motion quality, long-term coherence, and sampling efficiency.
@@ 1 Introduction
Figure 1: ReactDance generates high-fidelity, long-form reactive dance. Conditioned on a leader and music, our model uses a hierarchical representation to refine motion from coarse (translucent) to fine-grained (solid). Its parallel sampling mechanism produces coherent sequences exceeding 2000 frames within 2 seconds, avoiding the quality degradation and slow speed of autoregressive sampling.
Reactive Dance Generation (RDG) aims to synthesize lifelike and artistically coherent reactor motions that aligns with a lead dancer and accompanying music. This form of responsive interaction embodies a complex interplay of artistic expression and interpersonal dynamics that is a key challenge in fields like social robotics and human-agent interaction. The ability to generate such intricate, reactive behaviors holds significant potential for applications like virtual avatar animation for gaming and the metaverse (Hu, 2024; Zhang et al., 2023), as well as embodied AI and human-robot interaction (Granados et al., 2017; Chen et al., 2025).
Despite its promise, RDG faces two critical, unaddressed challenges: (1) modeling fine-grained spatial interactions and (2) maintaining long-term temporal coherence. While recent methods have advanced duet synchronization using tailored network architectures (Yao et al., 2023; Liang et al., 2024; Ghosh et al., 2024), physical constraints (Wang et al., 2024; Xu et al., 2024), or reinforcement learning (Siyao et al., 2024), they fail to resolve these core issues and fall short of generating truly authentic dance sequences. Their reliance on holistic, high-level constraints overlooks the subtle yet decisive local movements (e.g., the whip-like ’boleo’ in Tango), producing movements that are synchronized yet artistically sterile. Moreover, a fundamental difficulty for RDG is ensuring long-term temporal coherence. Models are typically trained on short clips, a practice driven by both the algorithmic complexity of learning long-term dependencies and practical computational constraints. This creates an inherent discrepancy between training and inference, leading to error accumulation that manifests as temporal drift and an eventual breakdown of synchronization.
To address these limitations, we draw inspiration from structural principles of dance theory and propose two core concepts for generating high-fidelity reactive motion:(1) Hierarchical Motion Composition: The principle that dance performance is inherently layered, with foundational full-body rhythms providing a coarse structure and fine-grained limb movements adding detailed semantics (Seham, 2016).(2) Modular Temporal Coherence: The observation that long-form choreography achieves consistency by composing shorter, coherent motion phrases and ensuring smooth transitions between them (Krug, 2022).
@@ 2.1 Music-Driven Dance Generation
Music-driven dance generation has advanced from solo to multi-dancer interactions. Early music-driven dance generation relied on sequence-to-sequence models like LSTMs (Tang et al., 2018) and GANs (Li et al., 2022), but often suffered from rigid transitions and a lack of detail. While auto-regressive models (Li et al., 2021; Siyao et al., 2022) improved solo dance quality, they lacked mechanisms for duet interactions. Recent duet-specific models still fall short; approaches like DanY (Yao et al., 2023) and Duolando (Siyao et al., 2024) use single-scale representations, causing synchronization artifacts, while DuetGen’s (Ghosh et al., 2025) coupled representation limits motion diversity. In contrast, we introduce a hierarchical representation that captures multi-scale spatial dynamics to generate high-fidelity, coherent reactive duets.
@@ 2.2 Hierarchical Motion Generation Models
The quality and controllability of generated motion are critically dependent on its underlying representation. Codebook-based autoregressive models (Gong et al., 2023; Jiang et al., 2024) quantize motion into discrete tokens but are prone to error accumulation and detail loss from VQ-VAE (Oord et al., 2017) quantization. While diffusion models (Qi et al., 2025; Liu et al., 2025) achieve state-of-the-art fidelity, they typically operate on flat representations, lacking a structure for multi-level control. Existing hierarchical models (Qi et al., 2023; Li et al., 2024) have focused on temporal, not spatial, refinement. Our work fills this gap by introducing a spatial hierarchy that enables both high-fidelity synthesis and multi-scale control over the generated motion.
@@ 3.1 Motion and Music Representation
The leader’s motion is modeled holistically as a single sequence of global joint positions, denoted as $\textbf{M}_{L}\in\mathbb{R}^{N\times J\times 3}$. In contrast, the reactor’s motion is decomposed into three independent components: (1) Upper-body motion ($\textbf{M}_{Rup}\in\mathbb{R}^{N\times J_{up}\times 3}$): The localized positions of the $J_{up}$ upper-body joints relative to the reactor’s own root joint. (2) Lower-body motion ($\textbf{M}_{Rdown}\in\mathbb{R}^{N\times J_{down}\times 3}$): The localized positions of the $J_{down}$ lower-body joints relative to the reactor’s own root joint. (3) Relative root distance ($\textbf{M}_{tr}\in\mathbb{R}^{N\times 3}$): The displacement vector between the leader’s and reactor’s roots.
Music Representation. For the accompanying music, we use Librosa (Brian McFee et al., 2015) to extract a rich acoustic feature vector $\bm{c}\in\mathbb{R}^{54}$ for each corresponding motion frame. This vector comprises 20-dimensional MFCCs, their 20-dimensional deltas, 12-dimensional chroma features, and 1-dimensional onset strength and beat tracking signals.
Our goal is thus to generate the set of reactor motion components $\textbf{M}_{R}=\{\textbf{M}_{Rup},\textbf{M}_{Rdown},\textbf{M}_{tr}\}$ conditioned on the music $\bm{c}$ and the leader’s motion $\textbf{M}_{L}$. To process the complete reactor motion $\textbf{M}_{R}$, each of its components is passed through a dedicated and structurally identical HFSQ stream. For clarity, the following sections will describe the core methodology for a single holistic stream.
@@ 3.2 Hierarchical Motion Representation via HFSQ
Encoder and Quantizer. A temporal encoder, consisting of 1D convolutional layers, maps an input reactor motion sequence $\textbf{M}_{R}\in\mathbb{R}^{N\times J_{R}}$ to a downsampled feature sequence $\bm{v}\in\mathbb{R}^{n\times d}$. This sequence is then processed by our HFSQ module, which operates in two stages:
@@
As we will detail, this structured set $\mathcal{V}$ serves as the target space for our diffusion model. Notably, we find that a lightweight configuration with just two residual quantization layers (R=2) is sufficient for our final architecture.
@@
$\mathcal{L}_{\text{kinematic}}$ $=\sum_{k\in\{\text{pos},\text{vel},\text{acc}\}}\lambda_{Pk}\cdot\|\textbf{M}^{(k)}_{R}-\hat{\textbf{M}}^{(k)}_{R}\|_{1},$ (4)
@@
The $\textbf{M}^{(k)}$ denotes the $k$-th order temporal derivative (position, velocity, acceleration) of the ground-truth motion, and $\hat{\textbf{M}}^{(k)}$ denotes the reconstructed counterparts. This objective ensures that the final reconstruction is both physically realistic and faithful to the encoder’s output.
@@ 3.3 Latent Diffusion on Hierarchical Features
Diffusion Framework. We model the generation process using a DDPM (Ho et al., 2020). The forward process gradually adds Gaussian noise to the HFSQ latents $\bm{x}_{0}=\mathcal{V}$ over T timesteps. The reverse process trains a Transformer-based network $\mathcal{G}_{\theta}$ to denoise the corrupted latents $\bm{x}_{t}$ at timestep $t$, conditioned on music c and leader motion $\textbf{M}_{L}$. To leverage the hierarchical structure of $\mathcal{V}$, we apply independent weights to the loss for each residual scale. The latent diffusion loss is:
$\mathcal{L}_{\text{simple}}=\sum_{r=1}^{R}\lambda^{r}\cdot\mathbb{E}_{\bm{x}_{0},t,\epsilon}\left[\|\mathcal{V}^{r}-\mathcal{G}_{\theta}(\bm{x}_{t},t,\bm{c},\textbf{M}_{L})^{r}\|_{2}^{2}\right],$ (6)
@@
3.4 Efficient Long Sequence Generation
Generating long, coherent dance sequences is a critical challenge, as the desired inference length $K$ often far exceeds the model’s training window $W$. A common approach is to apply diffusion models autoregressively—sequentially generating and stitching together short clips. However, this method has significant drawbacks: (1) slow inference due to its serial nature, (2) error accumulation, where early mistakes degrade the quality of later parts, and (3) blending artifacts at the overlaps between clips. To overcome these issues, we introduce the Blockwise Local Context (BLC), a non-autoregressive sampling strategy that generates all parts of a long sequence in parallel while ensuring both local and global coherence. BLC’s design rests on two principles: intra-block consistency and inter-block continuity.
Intra-block consistency ensures that each block is generated with high fidelity. We achieve this by aligning the temporal context of each inference block with that of the training window, using periodic applications of causal masking and positional encoding. Specifically, we apply a periodic causal attention mask $\mathcal{M}$ across the entire sequence of length $K$. This mask is a block-diagonal matrix(Fig. 4) that enforces causality strictly within each non-overlapping block of size $W$. This structure prevents information from leaking between blocks, thereby halting error propagation. Complementing the mask, we introduce a phase-aligned periodic positional encoding $\mathcal{P}_{i}$ that resets temporal phase for each block:
$\mathcal{P}_{i}=\sin\left(\frac{\pi(i\bmod W)}{W}\right)\oplus\cos\left(\frac{\pi(i\bmod W)}{W}\right),$ (8)
where $i$ is the global temporal position. This strategy makes every block’s positional context appear identical to the fixed-length window the model saw during training, ensuring it generalizes effectively regardless of a block’s actual position in the long sequence.
Inter-block continuity ensures seamless transitions between these parallel blocks. This is achieved by training the entire framework with a dense sliding window approach, with $m\ll T$ where $m$ is the sliding stride. By exposing the model to a vast number of overlapping subsequences during training, it implicitly learns how to naturally begin and end motion phrases from any temporal point. During inference, this learned behavior ensures that the overlapping regions between blocks are generated with smooth, natural transitions, creating a coherent long-form performance without costly sequential generation or explicit blending.
@@ 3.5 Layer-Decoupled Classifier-Free Guidance
Classifier-free guidance (CFG) is a powerful technique for controlling the trade-off between fidelity and diversity in generative models (Ho & Salimans, 2021). However, in motion generation, conventional CFG applies a single guidance weight across the entire motion representation (Tevet et al., 2023; Shafir et al., 2024b). This is a significant limitation for complex, hierarchical tasks like reactive dance, as it cannot distinguish between different motion semantics encoded at different scales (e.g., overall body posture vs. fine-grained limb articulation).
To address this, we propose Layer-Decoupled Classifier-Free Guidance (LDCFG), a method that enables fine-grained, multi-scale control over the generation process. As shown in Fig. 4, LDCFG assigns an independent guidance strength to each scale of our HFSQ representation:
@@
where $s_{r}$ is the guidance strength for the $r$-th residual scale.

### v3

@@ Abstract
Reactive dance generation (RDG), the task of generating a dance conditioned on a leader’s motion and music, holds significant promise for enhancing human-robot interaction and immersive digital entertainment. Despite progress in duet synchronization and motion-music alignment, two key challenges remain: generating fine-grained spatial interactions and ensuring long-term temporal coherence. In this work, we introduce ReactDance, a diffusion framework that operates on a novel hierarchical latent space to address these spatiotemporal challenges in RDG. First, for high-fidelity spatial expression and fine-grained control, we propose Hierarchical Finite Scalar Quantization (HFSQ). This multi-scale motion representation effectively disentangles coarse body posture from high-frequency dynamics, enabling independent and detailed control over both aspects through a layered guidance mechanism. Second, to efficiently generate long sequences with high temporal coherence, we propose Blockwise Local Context (BLC), a non-autoregressive sampling strategy. Departing from slow, frame-by-frame generation, BLC partitions the sequence into blocks and synthesizes them in parallel via periodic causal masking and positional encodings. Coherence across these blocks is ensured by a dense sliding-window training approach that enriches the representation with local temporal context. Extensive experiments show that ReactDance substantially outperforms state-of-the-art methods in motion quality, long-term coherence, and sampling efficiency. Project page: https://ripemangobox.github.io/ReactDance.
00footnotetext: 🖂 Corresponding Author.
@@ 1 Introduction
Figure 1: ReactDance generates high-fidelity, long-form reactive dance Conditioned on a leader and music, our model uses a hierarchical representation to refine motion from coarse (translucent) to fine-grained (solid). Its parallel sampling mechanism produces coherent sequences exceeding 2000 frames within 2 seconds, avoiding the quality degradation and slow speed of autoregressive sampling.
Reactive Dance Generation (RDG) aims to synthesize lifelike and artistically coherent reactor motions that align with a lead dancer and accompanying music. This form of responsive interaction embodies a complex interplay of artistic expression and interpersonal dynamics that is a key challenge in fields like social robotics and human-agent interaction. The ability to generate such intricate, reactive behaviors holds significant potential for applications like virtual avatar animation for gaming and the metaverse (Hu, 2024; Zhang et al., 2023), as well as embodied AI and human-robot interaction (Granados et al., 2017; Chen et al., 2025).
Despite its promise, RDG faces two critical, unaddressed challenges: (1) modeling fine-grained spatial interactions and (2) maintaining long-term temporal coherence. While recent methods have advanced duet synchronization using tailored network architectures (Yao et al., 2023; Liang et al., 2024; Ghosh et al., 2024), physical constraints (Wang et al., 2024; Xu et al., 2024), or reinforcement learning (Siyao et al., 2024), they fail to resolve these core issues and fall short of generating truly authentic dance sequences. Their reliance on holistic, high-level constraints overlooks the subtle yet decisive local movements (e.g., the whip-like boleo in Tango), producing movements that are synchronized yet artistically sterile. Moreover, a fundamental difficulty for RDG is ensuring long-term temporal coherence. Models are typically trained on short clips, a practice driven by both the algorithmic complexity of learning long-term dependencies and practical computational constraints. This creates an inherent discrepancy between training and inference, leading to error accumulation that manifests as temporal drift and an eventual breakdown of synchronization.
To address these limitations, we draw inspiration from structural principles of dance theory and propose two core concepts for generating high-fidelity reactive motion:(1) Hierarchical Motion Composition: The principle that dance performance is inherently layered, with foundational full-body rhythms providing a coarse structure and fine-grained dynamics adding detailed semantics (Seham, 2016).(2) Modular Temporal Coherence: The observation that long-form choreography achieves consistency by composing shorter, coherent motion phrases and ensuring smooth transitions between them (Krug, 2022).
@@ 2.1 Music-Driven Dance Generation
Music-driven dance generation has advanced from solo to multi-dancer interactions. Early music-driven dance generation relied on sequence-to-sequence models like LSTMs (Tang et al., 2018) and GANs (Li et al., 2022), but often suffered from rigid transitions and a lack of detail. While auto-regressive models (Li et al., 2021; Siyao et al., 2022) improved solo dance quality, they lacked mechanisms for duet interactions. Recent duet-specific models still fall short; approaches like DanY (Yao et al., 2023) and Duolando (Siyao et al., 2024) use single-scale representations, causing synchronization artifacts, while DuetGen’s (Ghosh et al., 2025) coupled representation models the joint distribution, inherently limiting the conditional diversity of reactor responses given a fixed leader. In contrast, we introduce a hierarchical representation that captures multi-scale spatial dynamics to generate high-fidelity, coherent reactive duets.
@@ 2.2 Hierarchical Motion Generation Models
The quality and controllability of generated motion are critically dependent on its underlying representation. Codebook-based models (Gong et al., 2023; Jiang et al., 2024) are prone to codebook collapse and detail loss from VQ-VAE (Oord et al., 2017) quantization. Although MoMask (Guo et al., 2024) introduces residual structures, its reliance on RVQ-VAE and serial generation causes cumulative errors across scales. While diffusion models (Qi et al., 2025; Liu et al., 2025) achieve state-of-the-art fidelity, they typically operate on flat representations, lacking a structure for multi-level control. Existing hierarchical models (Qi et al., 2023; Li et al., 2024) have focused on temporal, not spatial, refinement. Our work fills this gap by introducing a spatial hierarchy that enables both high-fidelity synthesis and multi-scale control over the generated motion.
@@ 3.1 Motion and Music Representation
The leader’s motion is modeled holistically as a single sequence of global joint positions, denoted as $\mathbf{M}_{L}\in\mathbb{R}^{N\times J\times 3}$. In contrast, the reactor’s motion is decomposed into three independent components: (1) Upper-body motion ($\mathbf{M}_{Rup}\in\mathbb{R}^{N\times J_{up}\times 3}$): The localized positions of the $J_{up}$ upper-body joints relative to the reactor’s own root joint. (2) Lower-body motion ($\mathbf{M}_{Rdown}\in\mathbb{R}^{N\times J_{down}\times 3}$): The localized positions of the $J_{down}$ lower-body joints relative to the reactor’s own root joint. (3) Relative root distance ($\mathbf{M}_{tr}\in\mathbb{R}^{N\times 3}$): The displacement vector between the leader’s and reactor’s roots.
Music Representation. For the accompanying music, we use Librosa (McFee et al., 2015) to extract a rich acoustic feature vector $\bm{c}\in\mathbb{R}^{54}$ for each corresponding motion frame. This vector comprises 20-dimensional MFCCs, their 20-dimensional deltas, 12-dimensional chroma features, and 1-dimensional onset strength and beat tracking signals.
Our goal is thus to generate the set of reactor motion components $\mathbf{M}_{R}=\{\mathbf{M}_{Rup},\mathbf{M}_{Rdown},\mathbf{M}_{tr}\}$ conditioned on the music $\bm{c}$ and the leader’s motion $\mathbf{M}_{L}$. To process the complete reactor motion $\mathbf{M}_{R}$, each of its components is passed through a dedicated and structurally identical HFSQ stream. For clarity, the following sections will describe the core methodology for a single holistic stream.
@@ 3.2 Hierarchical Motion Representation via HFSQ
Encoder and Quantizer. A temporal encoder, consisting of 1D convolutional layers, maps an input reactor motion sequence $\mathbf{M}_{R}\in\mathbb{R}^{N\times J_{R}}$ to a downsampled feature sequence $\bm{v}\in\mathbb{R}^{n\times d}$, where $n=N/4$. This sequence is then processed by our HFSQ module, which operates in two stages:
@@
Frequency-based Semantic Decomposition. Crucially, this residual hierarchy naturally disentangles motion semantics based on signal energy. The base layer ($r=1$), minimizing the primary reconstruction error, captures the Coarse Motion (e.g., global posture, orientation, and low-frequency trajectories). Subsequent layers ($r\geq 2$), encoding the residual error, capture the Fine Motion (e.g., high-frequency local dynamics and subtle articulation details). As we will detail, this structured set $\mathcal{V}$ serves as the target space for our diffusion model. Notably, we find that a lightweight configuration with just two residual quantization layers ($R=2$) is sufficient for our final architecture, with more details available in Appendix A.6.
@@
$\mathcal{L}_{\text{kinematic}}$ $=\sum_{k\in\{\text{pos},\text{vel},\text{acc}\}}\lambda_{Pk}\cdot\|\mathbf{M}^{(k)}_{R}-\hat{\mathbf{M}}^{(k)}_{R}\|_{1},$ (4)
@@
The $\mathbf{M}^{(k)}$ denotes the $k$-th order temporal derivative (position, velocity, acceleration) of the ground-truth motion, and $\hat{\mathbf{M}}^{(k)}$ denotes the reconstructed counterparts. This objective ensures that the final reconstruction is both physically realistic and faithful to the encoder’s output.
@@ 3.3 Latent Diffusion on Hierarchical Features
Diffusion Framework. We model the generation process using a DDPM (Ho et al., 2020). The forward process gradually adds Gaussian noise to the HFSQ latents $\bm{x}_{0}=\mathcal{V}$ over T timesteps. The reverse process trains a Transformer-based network $\mathcal{G}_{\theta}$ to denoise the corrupted latents $\bm{x}_{t}$ at timestep $t$, conditioned on music c and leader motion $\mathbf{M}_{L}$. To leverage the hierarchical structure of $\mathcal{V}$, we apply independent weights to the loss for each residual scale. The latent diffusion loss is:
$\mathcal{L}_{\text{simple}}=\sum_{r=1}^{R}\lambda^{r}\cdot\mathbb{E}_{\bm{x}_{0},t,\epsilon}\left[\|\mathcal{V}^{r}-\mathcal{G}_{\theta}(\bm{x}_{t},t,\bm{c},\mathbf{M}_{L})^{r}\|_{2}^{2}\right],$ (6)
@@
3.4 Efficient Long Sequence Generation via Blockwise Local Context
Generating long, coherent dance sequences is a critical challenge, as the desired inference length $K$ often far exceeds the model’s training window $T$. While autoregressive methods can extend sequence length, they suffer from serial inference latency and error accumulation. To overcome these limitations, we introduce Blockwise Local Context (BLC), a non-autoregressive sampling strategy that generates partitioned long sequences in parallel. BLC achieves high-fidelity generation through two synergistic principles: Intra-block Consistency (via sampling protocols) and Inter-block Continuity (via training dynamics).
Intra-block Consistency via Parallel Sampling. To enable parallel generation, BLC aligns the inference context of an arbitrary long sequence with the fixed-length context seen during training. We achieve this through:
  • Periodic Causal Masking: A block-diagonal attention mask $\mathcal{M}$ is applied across the sequence to enforce causality strictly within non-overlapping blocks of size $T$, preventing information leakage and halting error propagation between blocks.
  • Phase-aligned Positional Encoding: A periodic positional encoding $\mathcal{P}_{i}$ is introduced to reset the temporal phase for each block:
    $\mathcal{P}_{i}=\sin\left(\frac{\pi(i\bmod T)}{T}\right)\oplus\cos\left(\frac{\pi(i\bmod T)}{T}\right),$ (8)
    where $i$ is the global temporal position. This ensures that every inference block acts as an independent, well-formed window identical to the training distribution.
Inter-block Continuity via Dense Sliding Window (DSW). A core challenge in parallel block sampling is ensuring seamless transitions at the boundaries between independently generated blocks. We resolve this not by explicit post-processing, but by enforcing a Latent Manifold Constraint during training via our Dense Sliding Window (DSW) strategy. We train the model with a stride $m$ significantly smaller than the window size ($m\ll T$, e.g., $m=4,T=240$). This design is crucial for two reasons:
  • Phase-Agnostic Transition Prior: By processing motion windows at every possible phase shift, the model observes any given motion frame $f_{t}$ in diverse temporal roles (as a start, middle, or end frame). Consequently, the decoder learns a phase-agnostic transition function, understanding that adjacent latents $\bm{z}_{t}$ and $\bm{z}_{t+1}$ must resolve to continuous motion regardless of where the window boundary cuts.
  • Decoder as Kinematic Smoother: During inference, even if the parallel attention blocks create slight boundary discontinuities in the latent space, the HFSQ decoder—having been trained on thousands of overlapping, phase-shifted windows—projects these boundary latents onto the valid continuous motion manifold. This implicitly stitches blocks $i$ and $i+1$ with smooth, physically plausible transitions.
In summary, BLC’s masking enables efficiency through parallelism, while DSW’s dense supervision ensures coherence by learning robust boundary invariance.
@@ 3.5 Layer-Decoupled Classifier-Free Guidance
Precision Control vs. Spatial Control. While our Multi-stream Input design (Section 3.1) explicitly handles Spatial Control by defining specific body regions for interaction, ReactDance requires a complementary mechanism for Precision Control—allowing users to modulate how strictly the generation adheres to conditioning signals across different semantic scales. Standard Classifier-Free Guidance (CFG) (Ho and Salimans, 2021) applies a single scalar weight across the entire motion representation (Tevet et al., 2023; Shafir et al., 2024b), forcing a coupled trade-off between global pose stability and fine-grained interaction details. This is suboptimal for hierarchical tasks where structural coherence and subtle dynamics may require different guidance strengths.
To address this, we propose Layer-Decoupled Classifier-Free Guidance (LDCFG). This method leverages the hierarchical nature of HFSQ for orthogonal control over motion structure and details.
Training Strategy: Independent Condition Dropout. To support decoupled guidance during inference, the model must learn to condition each hierarchical layer independently. Unlike standard CFG training which drops conditions for all layers simultaneously, we employ Independent Condition Dropout. During training, the conditioning signal ($\bm{c}$ and $\mathbf{M}_{L}$) for the $r$-th residual layer is replaced with a null embedding $\emptyset$ with an independent probability $p=0.2$. This forces the denoiser to learn the conditional distribution $P(\bm{z}^{r}|\bm{c},\mathbf{M}_{L})$ marginally, laying the foundation for layer-isolated control during inference.
Inference Mechanism. During sampling, LDCFG assigns an independent guidance strength $s_{r}$ to each scale of the HFSQ representation. The predicted noise is computed as:
@@
where $s_{r}$ controls the guidance scale for the $r$-th layer. This formulation provides an explicit mechanism for frequency-based control by modulating the coarse and fine components defined in Section 3.2. The Base Guidance ($s_{1}$): Modulates the Coarse Motion. Increasing $s_{1}$ anchors the global posture and orientation, ensuring kinematic stability. While the Residual Guidance ($s_{r\geq 2}$): Modulates the Fine Motion. Increasing $s_{2}$ specifically enhances the subtle local dynamics and interaction nuances without altering the macro-level motion structure.
This disentanglement allows ReactDance to act as a “Detail Specialist” or “Structure Specialist” as needed, achieving the optimal balance between fidelity and diversity.